# A Watermark for Low-entropy and Unbiased Generation in Large Language Models

### v2

@@ 5.4 Attacking STA
We assessed the robustness of different watermarking methods under different attacks consisting of the copy-paste attack (Kirchenbauer et al., 2023a), paraphrasing using GPT-3.5, and two configurations of the DIPPER attack (Krishna et al., 2024). Detailed settings of different attacks are described in Appendix F.4.
Table 4 reports the F1-score and AUC of watermark detection under each attack with $z=2$. As reported, for unbiased watermarks, RDW achieves the best result since its detection framework based on brute force search is designed to solve the robustness issue (Kuditipudi et al., 2023). In contrast, STA-M is robust against different attacks with a low detection time. For the copy-paste attack, since STA-M is based on the green-red list partition and changing a token can only affect the detection score of itself and the next token, it is naturally robust to simple text insertion and removal (Kirchenbauer et al., 2023a). Meanwhile, LLM-based attacks, such as GPT-3.5 and DIPPER, are designed to replace tokens in given texts by sampling from the LLM. STA-M effectively increases the proportion of green-list tokens by raising their probability in high-entropy scenarios without compromising too much text quality, making it difficult for LLM-based attacks to replace a substantial number of tokens in STA-M-generated text and remove the watermark. In conclusion, STA-M can generate text with high watermark strength against various attacks.
6 Conclusions and Future Work
In this work, we propose a new unbiased watermarking method named STA-1. We clarify the text quality (regarding the risk of unsatisfactory outputs under the same expectation) and watermark strength tradeoff of unbiased watermarks in low-entropy scenarios. We also extend STA-1 to STA-M which can enhance watermark strength with small text quality shifts. Experimental results on low-entropy datasets prove that STA-1 is comparable to other unbiased watermarks and has a low risk. Moreover, results from the high-entropy dataset demonstrate the efficiency of STA-1 and the robustness of STA-M. Future work of our study can be conducted in several ways. First, watermarking low-entropy tasks is still challenging and future work can devise better watermarking methods. Second, future work could incorporate more datasets and generative LLMs for evaluation of our method. Third, it is also possible to consider context code history to extend the unbiased results from the token level to the sequence level.
@@ 7 Related Work
With the development of LLMs, the idea of watermarking LLMs has been proposed (Aaronson, 2022; Kirchenbauer et al., 2023a) and widely explored. Existing white-box watermarking techniques can be categorized into watermarking during logits and probabilities generation (Kirchenbauer et al., 2023b; Lee et al., 2023; Hu et al., 2024; Wang et al., 2023; Fernandez et al., 2023; Zhao et al., 2023; Yoo et al., 2023; Ren et al., 2023; Takezawa et al., 2023), and watermarking by controlling sampling strategies (Christ et al., 2023; Kuditipudi et al., 2023; Hou et al., 2023; Fairoze et al., 2023). We refer readers to Appendix G for a detailed discussion on related work.
@@ Appendix A Research Gap Summary
Table 5: Research Gap Summary. Black-box refers to only black-box LLM needed during detection; Efficiency refers to efficient detection; Robustness refers to the robustness against simple watermarking attacks. Guarantee refers to the statistical guarantee of type II error; Unbiased refers to the same expectation requirement in low-entropy scenarios.
Literature Black-box Efficiency Robustness Guarantee Unbiased
Kirchenbauer et al. (2023a)
Hu et al. (2024)
Wu et al. (2024)
Christ et al. (2023)
Kuditipudi et al. (2023)
Fairoze et al. (2023)
STA-1
@@ Appendix B Details of Previous Methods
$P_{M,w}(x^{t}=j|x^{-N_{p}:(t-1)};k)=\frac{\exp\left(l_{j}^{t}+\mathds{1}_{\text{Green}}(j)\delta\right)}{\sum_{i\in\text{Red}}\exp(l_{i}^{t})+\sum_{i\in\text{Green}}\exp(l_{i}^{t}+\delta)},$
@@
$F(j|\widetilde{\mathcal{V}})=\max\left[\sum_{i\in\widetilde{\mathcal{V}}:i\leq j}P_{M}(x^{t}=i|\cdot)-\alpha,0\right]+\max\left[\sum_{i\in\widetilde{\mathcal{V}}:i\leq j}P_{M}(x^{t}=i|\cdot)-(1-\alpha),0\right].$
@@ C.1 Proof of Theorem 1
$h_{j}(p)=\mathbb{E}_{G,R:j\in G}\left[p_{j}^{w,k}\right]=\mathbb{E}_{G,R:j\in G}\left[p_{j}+\left(\sum_{i\in R}p_{i}\right)p_{j}\right].$
@@
$f_{j}(p)=\mathbb{E}_{G,R:j\in R}\left[p_{j}^{w,k}\right]=\mathbb{E}_{G,R:j\in R}\left[\left(\sum_{i\in R}p_{i}\right)p_{j}\right].$
@@
$=\gamma p_{j}+\gamma\frac{N_{R}}{N-1}(1-p_{j})p_{j}+(1-\gamma)p_{j}^{2}+(1-\gamma)\frac{(N_{R}-1)}{N-1}(1-p_{j})p_{j}$
$=\left(\gamma+\frac{N_{R}-(1-\gamma)}{N-1}\right)p_{j}+\left((1-\gamma)-\frac{N_{R}-(1-\gamma)}{N-1}\right)p_{j}^{2}$
@@ C.2 Proof of Theorem 2
$=p_{j}\left[\frac{N-1+N_{R}(1-p_{j})}{N-1}\right]$
$\geq p_{j}\left[\frac{N+N_{R}(1-p_{j})}{N}\right]$
$=p_{j}+(1-\gamma)p_{j}(1-p_{j}),$
where the inequality holds because the denominator is less than the numerator, and adding 1 to both leads to a decrease in the value.
$\mathbb{P}(G)\coloneqq\mathbb{P}(\text{sampling a token}\in G)$ $=\sum_{j=1}^{N}\gamma\mathbb{E}_{G,R:j\in G}\left[p_{j}^{w,k}\right]$
$\geq\gamma\sum_{j=1}^{N}p_{j}+(1-\gamma)p_{j}(1-p_{j})$
$=\gamma+\gamma(1-\gamma)\sum_{j=1}^{N}p_{j}(1-p_{j}).$
$\mathbb{P}(G^{t})\geq\gamma+\gamma(1-\gamma)\sum_{j=1}^{N}p_{j}^{t}(1-p_{j}^{t})=\gamma+\gamma(1-\gamma)Gini(p^{t}).$
$\mathbb{E}(|S|_{G})$ $=T\mathbb{E}_{t}\left[\mathbb{P}(G^{t})\right]\geq T\mathbb{E}_{t}\left[\gamma+\gamma(1-\gamma)Gini(p^{t})\right]$
$\geq T\left[\gamma+\gamma(1-\gamma)Gini^{*}\right]=\gamma T+(1-\gamma)\gamma TGini^{*},$
$\leq T\left[\gamma+(1-\gamma)\gamma Gini^{*}\right]\left[1-\gamma-(1-\gamma)\gamma Gini^{*}\right],$
C.3 Proof of Corollary 1
$=P(|S|_{G}-\mathbb{E}(|S|_{G})\leq\gamma T+\tilde{z}\sqrt{\gamma(1-\gamma)T}-\mathbb{E}(|S|_{G})|H_{a})$
$\leq P(|S|_{G}-\mathbb{E}(|S|_{G})\leq\gamma T+\tilde{z}\sqrt{\gamma(1-\gamma)T}-\underline{\mathbb{E}}|H_{a})$
[…]
$\leq\frac{\overline{\mathbb{V}}}{\overline{\mathbb{V}}+(\underline{\mathbb{E}}-\gamma T-\tilde{z}\sqrt{\gamma(1-\gamma)T})^{2}},$
where Cantelli’s inequality holds because
$\underline{\mathbb{E}}-(\gamma T+\tilde{z}\sqrt{\gamma(1-\gamma)T})=\gamma(1-\gamma)TGini^{*}-\tilde{z}\sqrt{\gamma(1-\gamma)T}>0$
C.4 Proof of Theorem 3
@@
We start with the variance calculation for the STA-1 method. Because STA-1 is an unbiased watermark by Theorem 1, we have $\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]=\mathbb{E}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$. Considering the identical uniformly random partition process of green and red lists associated with the uniformly distributed key $k$ as in the proof of Theorem 1, depending on whether the token $x_{max}$ is assigned to the green list or not initially, $p_{max}^{w,k}$ have two possible realizations:
@@
$\mathbb{E}_{G,R:x_{max}\in G}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]=\left[\frac{N_{R}(1-p_{max})p_{max}}{(N-1)}\right]^{2}.$
@@
$\mathbb{E}_{G,R:x_{max}\in R}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]=\left[\left(\frac{(N_{R}-1)(1-p_{max})}{N-1}+p_{max}\right)p_{max}-p_{max}\right]^{2}.$
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]=\mathbb{E}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$=\gamma\mathbb{E}_{G,R:x_{max}\in G}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]+(1-\gamma)\mathbb{E}_{G,R:x_{max}\in R}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$=\gamma\left[\frac{N_{R}(1-p_{max})p_{max}}{(N-1)}\right]^{2}+(1-\gamma)\left[\left(\frac{(N_{R}-1)(1-p_{max})}{N-1}+p_{max}\right)p_{max}-p_{max}\right]^{2}$
$=p_{max}^{2}(1-p_{max})^{2}\left[\gamma\frac{N_{R}^{2}}{(N-1)^{2}}+(1-\gamma)\frac{N_{G}^{2}}{(N-1)^{2}}\right]$
@@
Next, we compute the variance for the Dipmark method with a partition hyperparameter $\alpha$. Note that $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{E}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$ holds because Dipmark is also unbiased. In Dipmark, the uniformly distributed key $k$ controls the randomness of permutations. Under the same assumption that $p_{i}=(1-p_{max})/(N-1)$ for $i\neq x_{max}$, the relative orders among these $(N-1)$ tokens become irrelevant in the permutation. Therefore, there are a total of $N$ unique permutations, each with a probability of $1/N$. Specifically, in the first unique permutation, there are 0 tokens $i$ where $i\neq x_{max}$ placed to the left of $x_{max}$ and $(N-1)$ tokens $i$ where $i\neq x_{max}$ placed to the right of $x_{max}$. In the second one, there is 1 token on the left and $(N-2)$ tokens on the right, and so forth. The last permutation has $(N-1)$ tokens on the left and 0 on the right. If $j$ such tokens are on the left of $x_{max}$, $j=0,1,\cdots,(N-1)$, the corresponding $p_{max}^{w,k}$ is
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]$ $=\mathbb{E}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$=\frac{1}{N}\sum_{j=0}^{N-1}\left[p_{max}-1+2j\frac{(1-p_{max})}{(N-1)}\right]^{2}$
$=-(p_{max}-1)^{2}+\frac{1}{N}\sum_{j=0}^{N-1}4j^{2}\frac{(1-p_{max})^{2}}{(N-1)^{2}}$
$=(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}.$
Note that, this variance value does not depend on $\alpha$. When $\alpha=0.5$, Dipmark becomes $\gamma$-reweight. Therefore, $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]$.
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right]$ $=\mathbb{E}_{\Pi}\left(\mathbb{V}_{u}^{\text{RDW}}\left[p_{max}^{w,k}|\Pi\right]\right)+\mathbb{V}_{\Pi}\left(\mathbb{E}_{u}^{\text{RDW}}\left[p_{max}^{w,k}|\Pi\right]\right)$
@@
To compare $\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]$ and $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]$, consider
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]$ $=p_{max}^{2}(1-p_{max})^{2}\gamma(1-\gamma)\frac{N^{2}}{(N-1)^{2}}$
$<\frac{1}{4}(1-p_{max})^{2}\frac{N^{2}}{(N-1)^{2}}$
$=(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}\times\frac{3}{4}\frac{N^{2}}{N^{2}-1},$
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}=\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right].$
@@
$\leq p_{max}(1-p_{max})=\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right],$
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right],$

### v3

@@ 4.4 Attacking STA
Figure 3: Attacking Watermarks for C4. For baselines, we report the highest AUC score of unbiased and biased watermarks against each attack. Full results are available in Appendix F.4 and Table 6.
We assess the robustness of different watermarking methods under various attacks, including the copy-paste attack (Kirchenbauer et al., 2023a), paraphrasing using GPT-3.5, and two configurations of the DIPPER attack (Krishna et al., 2024). We plot the AUC of watermark detection with $z=2$ for STA-1 and STA-16, alongside the highest AUC values of biased and unbiased benchmarks against each attack in Figure 3. As depicted, on the one hand, the unbiased STA-1 method achieves satisfactory performance, matching the best-performing unbiased benchmark in each attack. This empirically demonstrates that the STA-1 is also robust to various attacks. On the other hand, by repeating sampling as high-entropy steps, the STA-16 method achieves better robustness than KGW. Detailed attack settings, results, and analysis can be found in Appendix F.4.
@@ 5 Related Work
With the development of LLMs, the idea of watermarking LLMs has been proposed (Aaronson, 2022) and widely explored (Tu et al., 2024). Existing white-box watermarking techniques can be categorized into watermarking during logits and probabilities generation (Wang et al., 2023; Zhao et al., 2023; Yoo et al., 2023; Ren et al., 2023; Takezawa et al., 2023; Lu et al., 2024), and watermarking by controlling sampling strategies (Christ et al., 2023; Kuditipudi et al., 2023; Hou et al., 2023; Fairoze et al., 2023). We refer readers to Appendix G for a detailed related work.
6 Conclusions
In this work, we propose a novel watermarking method named STA-1. Theoretically, we show that STA-1 is unbiased and has a lower risk than existing unbiased watermarks. During detection, STA-1 also provides statistical test guarantees on the type II error of watermark detection, demonstrates high efficiency in detection time, and remains robust against various watermarking attacks. Experimental results on public datasets show that STA-1 achieves the above properties simultaneously. We also extend STA-1 to STA-M, which can enhance watermark strength with small text quality shifts.
7 Limitations
We acknowledge several limitations in this work and suggest directions accordingly for future improvement. First, watermarking low-entropy tasks remains challenging, and future work could devise better watermarking methods to improve watermark strength while maintaining text quality. Second, LLM watermarks should be robust against paraphrasing attacks like GPT-3.5 and DIPPER even in low-entropy scenarios. Future work can consider extending watermarking methods by enhancing robustness in these scenarios. Third, it may also be useful to consider context code history to extend the unbiased results from the token level to the sequence level.
@@ Appendix B Details of Previous Methods
$P_{M,w}(x^{t}=j|x^{-N_{p}:(t-1)};k)$
$\quad=\frac{\exp\left(l_{j}^{t}+\mathds{1}_{\text{Green}}(j)\delta\right)}{\sum_{i\in\text{Red}}\exp(l_{i}^{t})+\sum_{i\in\text{Green}}\exp(l_{i}^{t}+\delta)},$
@@
$F(j|\widetilde{\mathcal{V}})=\max\left[\sum_{i\in\widetilde{\mathcal{V}}:i\leq j}P_{M}(x^{t}=i|\cdot)-\alpha,0\right]$
$\quad+\max\left[\sum_{i\in\widetilde{\mathcal{V}}:i\leq j}P_{M}(x^{t}=i|\cdot)-(1-\alpha),0\right].$
Notably, Dipmark becomes $\gamma$-reweight (Hu et al., 2024) when $\alpha=0.5$.
@@ C.1 Proof of Theorem 1
$h_{j}(p)$ $=\mathbb{E}_{G,R:j\in G}\left[p_{j}^{w,k}\right]$
$=\mathbb{E}_{G,R:j\in G}\left[p_{j}+\left(\sum_{i\in R}p_{i}\right)p_{j}\right].$
@@
$f_{j}(p)$ $=\mathbb{E}_{G,R:j\in R}\left[p_{j}^{w,k}\right]$
$=\mathbb{E}_{G,R:j\in R}\left[\left(\sum_{i\in R}p_{i}\right)p_{j}\right].$
@@
$=\gamma p_{j}+\gamma\frac{N_{R}}{N-1}(1-p_{j})p_{j}$
$\quad+(1-\gamma)p_{j}^{2}+(1-\gamma)\frac{(N_{R}-1)}{N-1}(1-p_{j})p_{j}$
$=\left(\gamma+\frac{N_{R}-(1-\gamma)}{N-1}\right)p_{j}$
$\quad+\left((1-\gamma)-\frac{N_{R}-(1-\gamma)}{N-1}\right)p_{j}^{2}$
@@ C.2 Proof of Theorem 2
We start with the variance for the STA-1 method. Because STA-1 is an unbiased watermark by Theorem 1, we have $\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]=\mathbb{E}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$. Considering the identical uniformly random partition process of green and red lists associated with the uniformly distributed key $k$ as in the proof of Theorem 1, depending on whether the token $x_{max}$ is assigned to the green list or not initially, $p_{max}^{w,k}$ have two possible realizations:
@@
$\mathbb{E}_{G,R:x_{max}\in G}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$\quad=\left[\frac{N_{R}(1-p_{max})p_{max}}{(N-1)}\right]^{2}.$
@@
$\mathbb{E}_{G,R:x_{max}\in R}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]=$
$[(\frac{(N_{R}-1)(1-p_{max})}{N-1}+p_{max})p_{max}-p_{max}]^{2}.$
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]$
$=\mathbb{E}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$=\gamma\mathbb{E}_{G,R:x_{max}\in G}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$\quad+(1-\gamma)\mathbb{E}_{G,R:x_{max}\in R}^{\text{STA-1}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$=p_{max}^{2}(1-p_{max})^{2}$
$\quad\left[\gamma\frac{N_{R}^{2}}{(N-1)^{2}}+(1-\gamma)\frac{N_{G}^{2}}{(N-1)^{2}}\right]$
@@
Next, we compute the variance of Dipmark with a partition hyperparameter $\alpha$. $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{E}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$ holds because Dipmark is also unbiased. In Dipmark, the uniformly distributed key $k$ controls the randomness of permutations. Under the same assumption that $p_{i}=(1-p_{max})/(N-1)$ for $i\neq x_{max}$, the relative orders among these $(N-1)$ tokens become irrelevant in the permutation. Therefore, there are a total of $N$ unique permutations, each with a probability of $1/N$. Specifically, in the first unique permutation, there are 0 tokens $i$ where $i\neq x_{max}$ placed to the left of $x_{max}$ and $(N-1)$ tokens $i$ where $i\neq x_{max}$ placed to the right of $x_{max}$. In the second one, there is 1 token on the left and $(N-2)$ tokens on the right, and so forth. The last permutation has $(N-1)$ tokens on the left and 0 on the right. If $j$ such tokens are on the left of $x_{max}$, $j=0,1,\cdots,(N-1)$, the corresponding $p_{max}^{w,k}$ is
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]$
$\quad=\mathbb{E}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[(p_{max}^{w,k}-p_{max})^{2}\right]$
$\quad=\frac{1}{N}\sum_{j=0}^{N-1}\left[p_{max}-1+2j\frac{(1-p_{max})}{(N-1)}\right]^{2}$
$\quad=-(p_{max}-1)^{2}+\frac{1}{N}\sum_{j=0}^{N-1}4j^{2}\frac{(1-p_{max})^{2}}{(N-1)^{2}}$
$\quad=(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}.$
Note that, this variance value does not depend on $\alpha$. When $\alpha=0.5$, Dipmark becomes $\gamma$-reweight. Then, $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]$.
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right]$ $=\mathbb{E}_{\Pi}\left(\mathbb{V}_{u}^{\text{RDW}}\left[p_{max}^{w,k}|\Pi\right]\right)$
$\quad+\mathbb{V}_{\Pi}\left(\mathbb{E}_{u}^{\text{RDW}}\left[p_{max}^{w,k}|\Pi\right]\right)$
@@
For the comparison between $\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]$ and $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]$, consider
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]$
$\quad=p_{max}^{2}(1-p_{max})^{2}\gamma(1-\gamma)\frac{N^{2}}{(N-1)^{2}}$
$\quad<\frac{1}{4}(1-p_{max})^{2}\frac{N^{2}}{(N-1)^{2}}$
$\quad=(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}\times\frac{3}{4}\frac{N^{2}}{N^{2}-1},$
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]$ $<(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}$
$=\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right].$
@@
$\leq p_{max}(1-p_{max})$
$=\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right],$
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]$ $<\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]$
$=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]$
$<\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right],$
@@
C.3 Proof of Theorem 3
$\geq p_{j}+\frac{N_{R}}{N}(1-p_{j})p_{j}$
$=p_{j}+(1-\gamma)p_{j}(1-p_{j}),$
where the inequality holds straightforwardly.
$\mathbb{P}(G)$ $\coloneqq\mathbb{P}(\text{sampling a token}\in G)$
$\quad=\sum_{j=1}^{N}\gamma\mathbb{E}_{G,R:j\in G}\left[p_{j}^{w,k}\right]$
$\quad\geq\gamma\sum_{j=1}^{N}p_{j}+(1-\gamma)p_{j}(1-p_{j})$
$\quad=\gamma+\gamma(1-\gamma)\sum_{j=1}^{N}p_{j}(1-p_{j}).$
$\mathbb{P}(G^{t})$ $\geq\gamma+\gamma(1-\gamma)\sum_{j=1}^{N}p_{j}^{t}(1-p_{j}^{t})$
$=\gamma+\gamma(1-\gamma)Gini(p^{t}).$
$\mathbb{E}(|S|_{G})$ $=T\mathbb{E}_{t}\left[\mathbb{P}(G^{t})\right]$
$\geq T\mathbb{E}_{t}\left[\gamma+\gamma(1-\gamma)Gini(p^{t})\right]$
$\geq T\left[\gamma+\gamma(1-\gamma)Gini^{*}\right]$
$=\gamma T+(1-\gamma)\gamma TGini^{*},$
$\leq T\left[\gamma+(1-\gamma)\gamma Gini^{*}\right]$
$\qquad\left[1-\gamma-(1-\gamma)\gamma Gini^{*}\right],$
C.4 Proof of Corollary 1
$\quad=P(|S|_{G}-\mathbb{E}(|S|_{G})\leq$
$\qquad\gamma T+\tilde{z}\sqrt{\gamma(1-\gamma)T}-\mathbb{E}(|S|_{G})|H_{a})$
$\quad\leq P(|S|_{G}-\mathbb{E}(|S|_{G})\leq$
$\qquad\gamma T+\tilde{z}\sqrt{\gamma(1-\gamma)T}-\underline{\mathbb{E}}|H_{a})$
$\quad\leq\frac{\mathbb{V}(|S|_{G})}{\mathbb{V}(|S|_{G})+(\underline{\mathbb{E}}-(\gamma T+\tilde{z}\sqrt{\gamma(1-\gamma)T}))^{2}}$
[…]
$\quad\leq\frac{\overline{\mathbb{V}}}{\overline{\mathbb{V}}+(\underline{\mathbb{E}}-\gamma T-\tilde{z}\sqrt{\gamma(1-\gamma)T})^{2}},$
where Cantelli’s inequality holds because
$\underline{\mathbb{E}}-(\gamma T+\tilde{z}\sqrt{\gamma(1-\gamma)T})$
$\quad=\gamma(1-\gamma)TGini^{*}-\tilde{z}\sqrt{\gamma(1-\gamma)T}>0$